OpenStudy (anonymous):

Can someone show me thoroughl steps on how tthis should be done please. An athlete executing a long jump leaves the ground at 28.0 angle and travels 7.80 m a)What was the take off speed? b)If this speed were increased by just 5.0%,how much longer would the jump be?

OpenStudy (anonymous):

@pkjag

OpenStudy (anonymous):

(a) Use \[v^{2}=u^{2}\sin^{2}\theta-2gs\] we know that \[v=0\] at maximum height, if we take 7.8m to be the vertical height \[s\], then substitute for \[\theta\] and \[s\] and find the take off speed \[u\] (b) Use the same equation here, just increase the answer you got from (a) by 5%, which should be \[u \times \frac{105}{100}\] then substitute that in our equation and solve for \[s\], then subtract 7.8 from your answer.

OpenStudy (anonymous):

for the 1st equation is it the same as vx=vox?

OpenStudy (anonymous):

it doesn't matter either way \[v=0\]

OpenStudy (anonymous):

ok thanks